# R-SFLLM: Jamming Resilient Framework for Split Federated Learning with Large Language Models

### v2

@@ IV-B2 Named Entity Recognition
Table II shows that BERT similarly achieves near-identical performance to the SFL baseline for CONLL2003 with F1 scores above $92\%$ for the scenarios Gaussian and Protection. Again, No Protection results in maximal disruption, with consistent F1 scores around $10\%$. This suggests that the model is either missing almost all of the entities (low recall) or predicting almost all entities incorrectly (low precision), which ultimately renders the obtained global model unusable for NER. Thus, worst-case jamming results in worst-case model performance. In addition, Figure 3g again shows that each client observes near optimal performance from early epochs on. This further corroborates the assumption that the BERT architecture is robust against noisy embeddings due to potentially unfair resource allocation in wireless SFL.
@@
RoBERTa achieves similar outcomes for CONLL2003 with near-identical performance to the original baseline with F1 scores of around $93\%$. Worst-case global model performance is observed for the scenario No Protection with F1 scores around $10\%$ as well. However, as observed for SC, the global SFL model for scenario Protection starts out at lower F1 scores around $60\%$ and converges gradually to the baseline toward the end of the global training as shown in Figure 3h. This observation can be attributed to the same argumentation as before, in which the RoBERTa base model seems to be more vulnerable to noise in word embeddings, especially if some clients are subjected to higher MSEs due to unfair resource allocation. Thus, similarly as for SC, the global SFL model performance converges to the original baseline after a few global rounds, where RoBERTa makes use of the adversarial training property and robustifies over the fine-tuning period.
In contrast to CONLL2003, WNUT17 represents a more challenging dataset, which not only exhibits a significantly lower baseline performance for non-distributed training, but also consists of only few samples, such that each client is challenged with poorer insights into the data. However, even in this case, BERT achieves good but not great performance for all relevant scenarios given the small dataset. Furthermore, the worst-case jammer is able to successfully impair the global model with F1 scores around $25\%$, which corresponds to the performance of the baseline after only one epoch as shown in Figure 3i. Accordingly, the global model in this case continues to miss almost all of the entities, having not progressed at all.
@@
Furthermore, RoBERTa’s sensitivity to noisy word embeddings becomes more pronounced when dealing with particularly small datasets. Specifically, Figure 5 shows that while clients 1 and 2 achieve the baseline performance, client 3 with higher MSE does not. This results in a worse performing global model as shown in Figure 3j, where the projected performance trajectory indicates that RoBERTa not only needs more data but also more global training rounds to successfully mitigate the worst-case jamming impact. However, the performance after 10 global rounds is already remarkably better than in the case where no R-SFLLM protection is applied. These results suggest that both BERT and RoBERTa become more susceptible to noisy embeddings in wireless SFL when small datasets with potentially non-representative samples are used for training. The model might thus not generalize well nor benefit from the adversarial training feature of R-SFLLM.
@@ IV-B3 Jamming only one User
Based on the previous discussions and the results in Figures 4 and 5, it suffices if only one SFL party performs suboptimal for the global model to decrease in performance. However, this was investigated under the setting that all SFL clients are jammed simultaneously and where always client 3 suffered from higher MSEs due to unfair resource allocation as a result of our water-filling approach. To extend the investigation, Figure 6 provides simulation results
@@
for BERT on SST2, where only one party (client 1) is jammed and shows that SFL is highly sensitive to aggregating even only one (randomly) poisoned model, regardless of whether BERT is more robust against noisy embeddings. Note that similar worst-case results can be obtained for RoBERTa, other datasets and other clients. This further corroborates the findings in [14], where it was shown that few poisoned clients already compromise the global model. We thus show that this can be achieved via jamming attacks in wireless SFL.
IV-B4 Barrage Jamming
Next, we quantify how adversaries other than worst-case jammers affect the SFL performance. To this end, we employ a barrage jammer with covariance $\boldsymbol{C}_{\boldsymbol{u}_{nk}}=P_{J}/(N_{J}NK)$ and demonstrate results for BERT and RoBERTa on CONLL2003 in Figures 7 and 8, respectively. In both cases, barrage jamming has a non-negligible impact on the performance for scenario No Protection, which at first resembles the worst-case but improves with later global rounds. For BERT, SFL performance gradually increases with every global round and achieves an F1 score of $51\%$, which is still $41\%$ less than the baseline. For RoBERTa, performance, increases only marginally after the sixth round and achieves an F1 score of $21\%$, being only $11\%$ higher than the worst-case. This is interesting as clients 1,2 and 3 observe MSEs of 5.3dB, 8.0dB and 9.2dB, respectively, which are significantly lower than for the worst-case but slightly higher than for scenario Protection. This indicates a nonlinear and architecture-dependent relationship where small noise acts as a regularizer but higher noise beyond a certain threshold results in severe corruption, particularly for noise-sensitive RoBERTa models. Nevertheless, only the worst-case jammer is able to consistently maintain its detrimental performance. As for scenario Protection, barrage jamming matches the previous experiments. Thus, R-SFLLM mitigates adversarial jamming with equal performance, regardless of whether worst-case, barrage, or consequently any other jammer is employed.
@@ V Conclusion
In this paper, we have investigated the problem of adversarial jamming attacks in wireless SFL with LLMs. We have demonstrated both theoretically and experimentally that jamming LLM word embeddings leads to significantly worse model outcomes. Specifically, we have shown that worst-case jamming results in worst-case performance across all SFL rounds, regardless of how many clients are being targeted. To address this, we have proposed R-SFLLM, a sensing-assisted anti-jamming framework that leverages information about the adversary’s DoAs to integrate resilience directly into the wireless system by design. We have justified this wireless approach to resilience by showing that the LLM model error is upper bounded by the communication MSE under a relaxed $(\boldsymbol{L}_{0},\boldsymbol{L}_{1})$ smoothness assumption, thereby emphasizing that the physical layer directly impacts the training. Extensive experiments validate the effectiveness of R-SFLLM and demonstrate the critical need to safeguard LLM word embeddings in SFL against jamming attacks to ensure model integrity in distributed training over wireless networks.
Appendix A Proof of Proposition 1
@@
[…]
Second, applying the norm |.||.|| . | and triangle inequality yields
@@
||| $L(\boldsymbol{y})-L(\boldsymbol{x})|\leq\left|\nabla L(\boldsymbol{x})^{T}(\boldsymbol{y}-\boldsymbol{x})\right|$
$+\left|\sum_{j=1}^{E}\left[L_{0,j}+L_{1,j}\left|\frac{\partial L(\boldsymbol{x})}{\partial x_{j}}\right|\right]|y_{j}-x_{j}|\cdot||\boldsymbol{y}-\boldsymbol{x}||_{2}\right|.$ (47)
By defining the vectors $\boldsymbol{u}=[u_{1}\ldots u_{E}]^{T}$ with $u_{j}=L_{0,j}+L_{1,j}\left|\frac{\partial L(\boldsymbol{x})}{\partial x_{j}}\right|$ and $\boldsymbol{v}=[v_{1}\ldots v_{E}]^{T}$ with $v_{j}=|y_{j}-x_{j}|$, above expression can be denoted via scalar products, i.e.
@@
The vector $\boldsymbol{u}$ can be further reformulated as $\boldsymbol{u}=\boldsymbol{L_{0}}+\boldsymbol{L_{1}}\odot|\nabla L(\boldsymbol{x})|$, where $\odot$ denotes the Hadamard product and $\nabla L(\boldsymbol{x})$ is the gradient vector of $L$. This eventually yields the statement in Lemma 2.
@@ Appendix B Proof of Proposition 2
With non-negative expectations, the Cauchy-Schwarz inequality, i.e. $\mathbb{E}[A\cdot B]\leq\sqrt{\mathbb{E}[A^{2}]\cdot\mathbb{E}[B^{2}]}$, can be applied. Further, as $\nabla_{\boldsymbol{e}_{q}}L(\boldsymbol{e}_{q})$ and $\boldsymbol{u}(\boldsymbol{e}_{q})$ are independent of $s_{qnk}$ and with $\mathbb{E}\left[\|\boldsymbol{e}_{q}-\hat{\boldsymbol{e}}_{q}\|_{2}^{2}\right]=\mathbb{E}\left[\|\boldsymbol{s}_{q}-\hat{\boldsymbol{s}}_{q}\|_{2}^{2}\right]$ from Proposition 1, we obtain the following statement of Proposition 2:

### v3

@@ IV-B2 Named Entity Recognition
Table II shows that BERT similarly achieves near-identical performance to the SFL baseline for the smaller CONLL2003 dataset with F1 scores above $92\%$ for the scenarios Gaussian and Protection. Again, No Protection results in maximal disruption, with consistent F1 scores around $10\%$. This suggests that the model is either missing almost all of the entities (low recall) or predicting almost all entities incorrectly (low precision), which ultimately renders the obtained global model unusable for NER. Thus, worst-case jamming results in worst-case model performance. In addition, Fig. 4g again shows that each client observes good performance from early epochs on. This further corroborates that the BERT architecture is robust against noisy embeddings, even for unfair resource allocation in wireless SFL setups.
@@
RoBERTa achieves similar outcomes for CONLL2003, matching the baseline performance with F1 scores of around $93\%$. Worst-case global model performance is observed for scenario No Protection with F1 scores around $10\%$ as well. However, as observed for SC, the global SFL model for scenario Protection starts out at lower F1 scores of around $60\%$ and converges gradually to the baseline toward the end of the global training, shown in Fig. 4h. This can be attributed to the same explanation as before, such that RoBERTa is more vulnerable to noisy embeddings, particularly when some clients are subjected to higher MSEs due to unfair resource allocation. Thus, similarly as for SC, the global SFL performance converges to the baseline after a few global rounds, where RoBERTa makes use of R-SFLLM’s adversarial training character and robustifies over the fine-tuning period.
@@
In contrast, WNUT17 represents a more challenging dataset, containing viral social media posts that are highly informal and complex to categorize for LLMs. Thus, baseline performance is significantly lower. In addition, WNUT17 consists of only few samples, such that each client is challenged with poorer insights into the data. However, even in this case, BERT achieves good but not great performance for all relevant scenarios given the small dataset. Again, the worst-case jammer is able to successfully impair the global model with F1 scores of around $25\%$, which corresponds to the performance of the baseline after only one epoch, shown in Fig. 4i. Hence, the global model continues to miss almost all of the entities, having not progressed at all. Furthermore, RoBERTa’s sensitivity to noisy embeddings due to unfair resource allocation becomes more pronounced when dealing with particularly small datasets, such as in this case. Specifically, Fig. 6 shows that while clients 1 and 2 attain the baseline, client 3 with higher MSE does not. This results in a worse performing global model, shown in Fig. 4j, where the projected performance trajectory indicates that RoBERTa not only needs more data but also more global SFL rounds to successfully mitigate the worst-case jamming impact. However, the performance after 10 global rounds is already significantly better than in the case without R-SFLLM protection. These results suggest that both BERT and RoBERTa become more susceptible to noisy embeddings in wireless SFL when small datasets with potentially non-representative samples are used for training. The model might thus not generalize well nor benefit from the adversarial training character of R-SFLLM.
@@ IV-B3 Jamming only one User
Based on the previous discussions and the results in Fig. 5 and Fig. 6, it suffices if only one SFL party performs suboptimal for the global model to decrease in performance. However, this was investigated under the setting that all SFL clients are jammed simultaneously and where always client 3 suffered from higher MSEs due to unfair resource allocation as a result of our water-filling approach. To extend the investigation, Fig. 7 provides simulation results for BERT on SST2, where only one party (client 1) is jammed, and shows that SFL is highly sensitive to aggregating even only one (randomly) poisoned model, regardless of whether BERT is more robust against noisy embeddings. Note that similar worst-case results can be obtained for RoBERTa, other datasets, and other clients. This further corroborates the findings in [14], where it was shown that few poisoned clients suffice to compromise the global model. Our results show that this can be achieved via jamming attacks in wireless SFL.
IV-B4 Barrage Jamming
Next, we quantify how adversaries other than worst-case jammers affect the SFL performance. To this end, we employ a barrage jammer with covariance $\boldsymbol{C}_{\boldsymbol{u}_{nk}}=P_{J}/(N_{J}NK)$ and demonstrate results for BERT and RoBERTa on CONLL2003 in Fig. 8 and Fig. 9, respectively. In both cases, barrage jamming has a non-negligible impact on the performance for scenario No Protection, which at first resembles the worst-case but improves with later global rounds. For BERT, SFL performance gradually increases with every global round and achieves an F1 score of $51\%$, which is still $41\%$ less than the baseline. For RoBERTa, performance increases only marginally after the sixth round and achieves an F1 score of $21\%$, being only $11\%$ higher than the worst-case. This is interesting as clients 1,2, and 3 observe MSEs of 5.3dB, 8.0dB, and 9.2dB, respectively, which are significantly lower than for the worst-case, but still higher than for scenario Protection. This shows that small noise acts as a regularizer but higher noise beyond a certain threshold results in severe corruption, particularly for noise-sensitive RoBERTa models.
@@ IV-B4 Barrage Jamming
Nevertheless, only the worst-case jammer is able to consistently maintain its detrimental performance. As for scenario Protection, barrage jamming matches the previous experiments. Thus, R-SFLLM mitigates adversarial jamming with consistent performance, regardless of whether worst-case, barrage, or any other jammer is employed, demonstrating its versatility across most practical jamming scenarios.
IV-C R-SFLLM Simulation Results for VLM Models.
Compared to BERT and RoBERTa, VLMs are multi-modal models that process both images and text in parallel. In particular, OpenAI’s CLIP ViT-B/16 VLM considered here uses a vision transformer (ViT) backbone, paired with a text transformer. Both encoders produce embeddings that are contrasted during training, i.e., matching image and text embeddings are pulled closer together while non-matching pairs are pushed apart, yielding multi-modal embeddings. This makes it particularly interesting to study the impact of adversarial noise on both language and image understanding. Furthermore, since a larger volume of corresponding embeddings must be transmitted during SFL, the system might become more susceptible to adversarial attacks, impacting resilience more severely than in lower-bandwidth NLP use cases. To investigate this, we fine-tune the CLIP VLM on eight diverse image classification datasets, including samples ranging from European satellite imagery in EuroSAT to German traffic signs in GTSRB. Results for all scenarios are shown in Table II, and Fig. 10 illustrates the model performance across global SFL rounds, averaged over all considered datasets.
In general, the SFL baseline converges after the third global round, achieving an average accuracy of 91%. This is partly due to variation in sample sizes, where smaller datasets, such as DTD and Cars, require more global rounds for convergence, while training on larger datasets like SVHN and MNIST typically converges after the first round. A similar trend is observed for scenarios Gaussian and Protection, both of which also converge by the third global round. While convergence is observed, the results clearly show the pronounced impact of noise for the more vulnerable multi-modal embeddings of the VLM architecture. Since both text and image embeddings are exposed to noise, the contrastive learning objective of VLMs is more strongly affected, leading to mismatching text labels for images, and thus to incorrect classification outcomes. This results in an initially larger performance gap between the SFL baseline and scenarios Gaussian and Protection. For example, after the first global round, the accuracy of Gaussian (Protection) is around 77% (74%), thereby 7% (10%) less than the SFL baseline. As training progresses, this performance gap narrows significantly, in particular, dropping to less than 1% after the third global round. These results highlight two key observations: First, the VLM’s multi-modal embeddings are indeed more susceptible to both benign (i.e., Gaussian) and adversarial noise during the early rounds of training. Second, R-SFLLM is able to maintain performance comparable to Gaussian, thus achieving close-to-baseline accuracies toward the end of training. This underscores R-SFLLM’s effectiveness in mitigating worst-case adversarial jamming across a variety of datasets, where No Protection would otherwise result in average performance as low as 8%, reflecting the model’s inability to correctly match images and text labels. Thus, across all classification tasks, R-SFLLM achieves performance close to the SFL baseline by the third global round, by which point the VLM has adapted to the adversarial training signal. Consistent with our previous NLP results, this holds even for the smallest datasets, such as DTD (4k samples) and Cars (8k samples), where R-SFLLM achieves accuracies within 2–3% of the baseline (see Table II). Furthermore, larger datasets such as MNIST remain highly competitive with 97.5% accuracy as more samples are observed during training. Thus, R-SFLLM enables the VLM to effectively learn contrastive embeddings even in the presence of adversarial jamming, resulting in strong performance across diverse datasets with varying sizes and distributions. This demonstrates that R-SFLLM is applicable across a range of transformer-based language models, including more complex multi-modal architectures.
Figure 10: Global SFL model performance for CV experiments with CLIP ViT-B/16, evaluated after each of the $N_{\text{rounds}}$ global SFL rounds and averaged across all datasets. R-SFLLM effectively safeguards multi-modal embeddings, achieving performance close to the SFL baseline after the third global round.
IV-D Ablation Results and Sensitivity Analysis.
We examine the impact of varying the number of users $Q$ and jamming antennas $N_{J}$ on the performance of R-SFLLM, demonstrating that our framework is both scalable and effective for various adversarial SFL setups.
IV-D1 Increasing the Number of Users $Q$
In Fig. 11, we compare sum rates for the scenarios No Protection, No Jammer, R-SFLLM, and Full Knowledge, where the latter represents an arbitrary, but optimal reference anti-jamming strategy that requires full knowledge about the jamming covariance matrix $\boldsymbol{C}_{\boldsymbol{u}_{nk}}$, as studied in [22]. We see that, in general, the sum rate increases with increasing number of users, and that our R-SFLLM framework approximates the sum rate for the optimal reference solution. In contrast, the worst-case jammer manages to bring down the sum rate to almost zero for scenario No Protection. This demonstrates that our proposed sensing-assisted anti-jamming strategy is highly competitive compared to established algorithms that require full system and jammer knowledge. To validate whether this translates directly into SFL performance, we fine-tune BERT on the larger QNLI binary classification dataset for SC with $Q=\{3,7,11,15\}$ users, respectively, and show corresponding performance plots in Fig. 12. In particular, we see that performance is sustained when increasing the number of SFL participants, such that outcomes only differ by negligible statistical variance around 1%. This demonstrates that R-SFLLM is highly scalable.
IV-D2 Increasing the Number of Jamming Antennas $N_{J}$
In Fig. 13, we compare sum rates across the same scenarios when equipping the adversarial jammer with more antennas $N_{J}$. We show that R-SFLLM maintains consistent performance, achieving rates close to the optimal reference solution and demonstrating that our proposed framework is resilient even against jammers with very large numbers of antennas. To extend this discussion, we further examined R-SFLLM’s anti-jamming performance across a wide range of jamming DoAs, power budgets, and scale parameters $\eta$ in our previous work in [22]. There, we showed that anti-jamming remains effective and scalable, independent of the number of wireless users or jammer capabilities, when $\eta$ is chosen to be significantly larger than the noise level $\sigma^{2}$, a practical heuristic that we carefully validated and discussed in more detail in Sec. III-B.
IV-E Summary of Results.
In our experiments, we showed that jamming LLM embeddings during SFL leads to severe performance degradation, with worst-case jamming resulting in worst-case training performance. Our investigation further revealed that some models are particularly sensitive to noisy embeddings. For example, RoBERTa is more susceptible due to the lack of stabilizing segment embeddings, while noise during CLIP VLM’s contrastive optimization challenges the learning of multi-modal embeddings. However, across all settings, our proposed R-SFLLM framework was able to effectively safeguard SFL training by mitigating noise from adversarial jammers, including worst-case ones, thereby ensuring resilience and achieving performance near identical to SFL baselines. To ensure wide-range applicability, we evaluated R-SFLLM for both NLP and CV domains using three different models (BERT, RoBERTa, and multi-modal CLIP ViT-B/16) across 13 different datasets of varying sizes and distributions, showing that R-SFLLM generalizes well to different types of transformer-based architectures in practice. In extensive ablation studies, we demonstrated that R-SFLLM is scalable and can be applied to large-scale SFL settings, while providing resilience against various types of adversarial jammers, regardless of their power, number of antennas, or system knowledge. Furthermore, we benchmarked R-SFLLM against an optimal reference solution, showing closely matching outcomes and thereby verifying its competitiveness against alternative anti-jamming strategies.
@@ V Conclusion
In this paper, we investigated the problem of adversarial jamming attacks in wireless SFL with language models. We demonstrated both theoretically and experimentally that jamming critical embedding parameters leads to significantly worse model outcomes. Specifically, we showed that worst-case jamming results in worst-case performance, regardless of how many clients are being targeted. To address this, we proposed R-SFLLM, a scalable, sensing-assisted anti-jamming framework that leverages information about the adversary’s DoAs to integrate resilience directly into the wireless system by design. We justified this wireless approach to resilience by showing that the LLM model error is upper bounded by the communication MSE under a relaxed $(\boldsymbol{L}_{0},\boldsymbol{L}_{1})$-smoothness assumption, thereby emphasizing that the physical layer directly impacts the training. Extensive experiments in both NLP and CV domains, using three different model architectures and 13 datasets, validate the effectiveness of R-SFLLM and demonstrate the critical need to safeguard embedding parameters in SFL against adversarial jamming attacks to ensure model integrity in distributed training over wireless networks.
Figure 11: Sum rate vs. number of users $Q$ for scenarios No Protection, No Jammer, R-SFLLM, and Full Knowledge. R-SFLLM approximates the optimal solution with Full Knowledge and increases the sum rate for increasing $Q$.
Figure 12: Global BERT SFL model performance for varying users $Q$. R-SFLLM remains effective and scalable when the number of SFL participants increases.
Figure 13: Sum rate vs. number of jamming antennas $N_{J}$ for scenarios No Protection, No Jammer, R-SFLLM, and Full Knowledge. R-SFLLM approximates the optimal reference solution and remains resilient against higher $N_{J}$.
Appendix A Proof of Lemma 2
@@
$+\int_{0}^{1}\langle\nabla L(\boldsymbol{x}+u(\boldsymbol{y}-\boldsymbol{x}))-\nabla L(\boldsymbol{x}),\boldsymbol{y}-\boldsymbol{x}\rangle du.$ (45)
Second, applying the norm $|.|$ and triangle inequality yields
$+\left|\int_{0}^{1}\langle\nabla L(\boldsymbol{x}+u(\boldsymbol{y}-\boldsymbol{x}))-\nabla L(\boldsymbol{x}),\boldsymbol{y}-\boldsymbol{x}\rangle du\right|.$ (46)
@@
By defining the vectors $\boldsymbol{u}=[u_{1}\ldots u_{E}]^{T}$ with $u_{j}=L_{0,j}+L_{1,j}\left|\frac{\partial L(\boldsymbol{x})}{\partial x_{j}}\right|$ and $\boldsymbol{v}=[v_{1}\ldots v_{E}]^{T}$ with $v_{j}=|y_{j}-x_{j}|$, above expression can be reformulated using scalar products, i.e.
@@
With $\boldsymbol{u}=\boldsymbol{L_{0}}+\boldsymbol{L_{1}}\odot|\nabla L(\boldsymbol{x})|$, where $\odot$ denotes the Hadamard product and $\nabla L(\boldsymbol{x})$ is the gradient vector of $L$, the statement in Lemma 2 is obtained.
@@ Appendix B Proof of Proposition 2
With non-negative expectations, the Cauchy-Schwarz inequality, i.e. $\mathbb{E}[A\cdot B]\leq\sqrt{\mathbb{E}[A^{2}]\cdot\mathbb{E}[B^{2}]}$, can be applied. Further, as $\nabla_{\boldsymbol{e}_{q}}L(\boldsymbol{e}_{q})$ and $\boldsymbol{u}(\boldsymbol{e}_{q})$ are independent of $s_{qnk}$, and with $\mathbb{E}\left[\|\boldsymbol{e}_{q}-\hat{\boldsymbol{e}}_{q}\|_{2}^{2}\right]=\mathbb{E}\left[\|\boldsymbol{s}_{q}-\hat{\boldsymbol{s}}_{q}\|_{2}^{2}\right]$ from Proposition 1, we obtain the statement of Proposition 2 as follows: